a solution of KIO3 is boiled with excess of oxalic cid. the products are K2C2O4,CO2,H2O&I2.CALCULATE quantity of oxalic acid in grams used in formation of one mole of iodine.


Grade:11

1 Answers

Yash Jain
9 years ago
The balanced eqn is given as
2KIO3 + 2(COOH)2 ---> K2C2O4 + 4CO2 + 2H2O + I2
             Oxalic acid
We observe that 2 moles of oxalic acid gives one mole of I2.
therefore, NOA = WOA / MOA = 2
So, req. mass of Oxalic acid (OA) = 2*MOA = 2*90 = 180g